8) CAL Electronic ware house sale AAA batteries, and is very popular item. Demand for this is normally distributed. Average demand was 40 packs per day and a standard deviation of 12. Lead time is constant and which is 4 days. A 95% service level was desired.

A) 436

B) 200

C) 295

D) 310

E) 356
Answer #1

DEMAND = 40
STDEV DEMAND = 12
SERVICE LEVEL = 95, CORRESPONDING Z VALUE 1.65
LEAD TIME = 4

REORDER POINT = (DEMAND * LT) + SAFETY STOCK
SAFETY STOCK = Z * STANDARD DEVIATION OF DEMAND * SQRT(LEAD TIME)

SAFETY STOCK = 1.65 * 12 * SQRT(4) = 39.36 OR 40

ROP = (40 * 4) + 40 = 200